OpenStudy (anonymous):

whats y = -3(x-2)(x+3)

jigglypuff314 (jigglypuff314):

you mean in standard form?

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

well then, it's the same concept as before standard form of parabola: y = ax^2 + bx + c so just foil and distribute

jigglypuff314 (jigglypuff314):

y = -3(x-2)(x+3) = -3(x^2+x-6) then distribute the negative 3

jigglypuff314 (jigglypuff314):

if you're not sure how to foil http://www.freemathhelp.com/using-foil.html

OpenStudy (anonymous):

i got y= -6 +21

jigglypuff314 (jigglypuff314):

-.- y = -3x^2 -3x + 18

OpenStudy (anonymous):

lol idk tf im doing